OpenStudy (anonymous):

Michel and Susan have a combined age of 17 years. Michel is 5 years older than Susan. How old is Michel. a. 5 b. 6 c. 11 d. 22

OpenStudy (anonymous):

Michel is 11 years old.

OpenStudy (texaschic101):

m + s = 17 s = m - 5 m + (m-5) = 17 2m - 5 = 17 2m = 17 + 5 2m = 22 m = 22/2 m = 11 so yes...michael is 11

OpenStudy (anonymous):

a snail is at the bottom of a 10 ft wall. Each day it climbs up 3 ft. each night it slides back down 1ft. on which day will it reach the top of the wall. a. day 3 b. day 4 c. day 5 d. day 6

OpenStudy (anonymous):

anybody here

OpenStudy (anonymous):

texaschic101 are you here

OpenStudy (anonymous):

the population of pinecrest is 18,000. lakewood's population is 21/25 of pinecrest's. what is lakewood's population. a. 21,149 b. 15,120 c. 741 d. 721

OpenStudy (anonymous):

help anybody

OpenStudy (anonymous):

b

OpenStudy (anonymous):

B And the best way to find the answer just use a calculator

OpenStudy (anonymous):

esther starts at one corner of a square field and walks 90 ft along one side to the another corner of the field. she turns 90 degreese and walks 70 feet, and then walks straight back to where she started. what is the area of the part of the field she walked around. a. 3150 ft2 b. 4900 ft2 c. 6300 ft2 d. 8100 ft2

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

anybody please help